OpenStudy (anonymous):

If f(x)=8x and g(x)= 2x+1, find (f og) (x).

OpenStudy (aum):

(f og)(x) = f(g(x)) f(x) = 8x f(g(x)) = 8g(x) = 8(2x+1) = 16x + 8

OpenStudy (anonymous):

16x^2*8x?

OpenStudy (sleepyhead314):

oh wait ^^ sorry I got confused by the notation @aum is correct

OpenStudy (anonymous):

wait im right?

OpenStudy (aum):

See my first reply.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

thank you

OpenStudy (aum):

yw.